# Semantic Edge Computing and Semantic Communications in 6G Networks: A Unifying Survey and Research Challenges

### v2

@@ 6 Conclusion
Semantic Edge Computing and Semantic Communication are two critical approaches to achieve collaborative intelligence for next-generation edge computing. However, there is no existing survey to systematically unify both fields and literature in one domain often neglect the progress in the other. In this work, we provided a unified overview and reviewed the recent breakthrough in both fields by addressing their technical contributions. In addition, we discussed the research challenge and potential direction for future endeavour. We hope this survey can inspire further research for collaborative intelligence.
@@ Acknowledgements
This work was supported by the U.S. Army Combat Capabilities Development Command (DEVCOM) Army Research Laboratory under Cooperative Sub-Award Agreement No. W911NF-24-2-0224.

### v3

@@ 6 Conclusion
Semantic Edge Computing and Semantic Communication are two critical approaches to achieve collaborative intelligence for next-generation edge computing. However, there is no existing survey to systematically unify both fields and literature in one domain often neglects the progress in the other. In this work, we provide a unified overview and review the recent breakthroughs in both fields by addressing their technical contributions. In addition, we discuss the research challenges and potential directions for future endeavors. We hope this survey can inspire further research for collaborative intelligence.
@@ Acknowledgements
This work was supported in part by the U.S. Army Combat Capabilities Development Command (DEVCOM) Army Research Laboratory under Cooperative Sub-Award Agreement No. W911NF-24-2-0224, by the National Science Foundation under grants CNS-2134973, ECCS-2229472, by the Air Force Office of Scientific Research under contract number FA9550-23-1-0261 and in part by the Office of Naval Research under award number N00014-23-1-2221.